# ECGrecover: a Deep Learning Approach for Electrocardiogram Signal Completion

### v1

@@ 4. Conclusions and future work
We tackled the problem of reconstructing the 12-lead ECG signal when only a partial subset of the signal is available. We proposed a U-Net-like model trained on a novel objective function, combining the classical mean squared error loss with the Pearson correlation. We extensively evaluated our method, ECGrecover, on a real-life ECG dataset comprising 15 119 ECGs. The comparison to recent state-of-the-art methods reveals the effectiveness of ECGrecover in recovering the missing signal, even in challenging cases such as the configuration, as quantified either with distortion metrics or by the preservation of QT segment or the QRS-complex.
The primary objective of this method is to enhance the comprehensive analysis of ECG data for clinicians. By utilizing the reconstruction capabilities of ECGrecover, we can significantly improve the readability of ECG traces, which can be determinant in diagnosing diseases. This advancement allows for a more detailed and accurate assessment of cardiac activity, facilitating better diagnostic decisions and patient care. Notably, the proposed approach is part of a larger translational research project, DeepECG4u555https://anr.fr/Project-ANR-20-CE17-0022, which aims to improve patient care through automatic analyses of ECG based on AI.
Appendix A Additional results
In this section, we discuss additional simulations and results that, due to space constraints, were not included in the main paper. For better readability, we have collected the comprehensive set of results in 19 tables, which can be found in Appendix B.
@@ A.1. On the parameter
In Section 2.2, we introduced the objective function in Equation 1 used to train our U-Net model
In the following section, we investigate the role of the hyperparameter to train the model. In particular, we consider , where for , and for , . We provide in Table 3 the results over the validation set of the five models trained with the different values of . As we can see, when the model is trained on the mean squared error alone, it focuses solely on minimizing the magnitude differences between the predicted and original ECG signals. Notably, while the results for RMSE and MAE remain consistent for the models with , the values for the PCC and DTW present higher variability, with an improvement of 4.65% in terms of PCC for the model with w.r.t. the model with and of 10.44% in terms of DTW. Interestingly, setting , resulted in the model performing well in terms of PCC but significantly worse across other metrics. This observation underscores the importance of employing a hybrid loss that considers the ECG signal’s spatial and temporal aspects.
For our testing phase, we chose based on its performance during validation, showing promising results compared to and , without extensive exploration of other values. However, a more detailed examination revealed to potentially offer better outcomes. This finding suggests that while provided satisfactory results, a further optimization of could enhance model performance. In future work, we plan to investigate this aspect more thoroughly to identify the most effective parameter value.
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | |
@@ A.2. Random masking
An illustration of an ECG subjected to such random masking, referred to as C, is depicted in Figure 7. For this process, a random start and end point between 0 and the total length of the signal were selected for each ECG lead. The segments outside the range between these two points were masked.
@@
The results, detailed in Table 4, indicate that ECGrecover consistently outperforms its competitors across various metrics, even in this challenging scenario. For instance, in the most difficult case concerning lead III, ECGrecover not only improves upon the best competitor, Pix2Pix, by 510% in terms of PCC but also surpasses it by approximately 77.78% in terms of DTW. Furthermore, ECGrecover demonstrates stable performance across different leads, contrasting with the variability seen in competitors like EKGAN and Pix2Pix, which show wide fluctuations in PCC and DTW scores.
A.3. ECGrecover on PTB-XL
@@
We present the evaluation of the proposed ECGrecover on a publicly available dataset, PTB-XL (Wagner et al., 2020). Specifically, the dataset contains 21799 recordings from 18869 patients of 10-second 12- lead ECGs sampled at 500Hz labeled as normal, myocardial infarction, ST/T change, conduction disturbance and hypertrophy. In Figure 10, we show the summary of the results. In particular, we followed the same experimental setting as the one already described for the Generepol dataset in Section 3.1.
To get an overall view, we also include the same representation of the results for the Generepol dataset in the figure.
A.4. Additional plots
@@
We include additional plots in Figure 8 and Figure 9, which could not be accommodated in the main paper due to space limitations. Specifically, in Figure 8, we illustrate a lead II signal reconstruction from input masks C3 and C5 across all methodologies evaluated in our study. Meanwhile, Figure 9 presents the ECG waveforms alongside their labeled wave coordinates (peaks) to enhance comprehension of the findings discussed in Section 3.2.2.
A.5. Experimental environment
The calculations were run on a Nvidia A100 80GB GPU, with the total training duration approximating 1 hour and 30 minutes.
Appendix B Tables
@@
This section includes the comprehensive numerical results of the evaluation. Specifically, Tables 5, 7 and 6 present the results for the PCC metric; Tables 8, 10 and 9 the results for RMSE metric; Tables 11, 13 and 12 the results for MAE metric; Tables 14, 16 and 15 the results for DTW metric.
@@
Finally, in Tables 17, 19 and 18 we include the results on the QT-segment preservation w.r.t. all the possible input masks. Additionally, we expanded the analysis on the QRS complex (see Figure 9) in Tables 20, 22 and 21. Moreover, in Tables 23, 25 and 24, we show the percentage of R peaks detected (%-R) in the reconstructed signals.
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |

### v3

@@ 4. Conclusions and future work
We tackled the problem of reconstructing the 12-lead ECG signal when only a partial subset of the signal is available. We proposed a U-Net-like model trained on a novel objective function, combining the mean squared error loss with the Pearson correlation. We evaluated ECGrecover on a real-life ECG dataset comprising 15 119 10-second ECGs. The comparison to recent SOTA revealed the effectiveness of ECGrecover in recovering the missing signal, even in challenging cases such as the C configuration, as quantified either with distortion metrics or by the preservation of the QT segment.
By using the reconstruction capabilities of ECGrecover, we can improve the readability of ECG traces, which can be determinant in diagnosing diseases. This advancement allows for a more detailed and accurate assessment of cardiac activity, facilitating better diagnostic decisions. Notably, ECGrecover is part of the DeepECG4U555https://anr.fr/Project-ANR-20-CE17-0022 translational research project, which aims to improve patient care through automatic analyses of ECG based on AI.
Acknowledgements
[…]
Appendix A Extended analysis of ECGrecover
This section presents additional analysis that could not be included in the main paper due to space constraints.
@@ A.1. On the parameter
In Section 2.2, we introduced the objective function in Equation 1 used to train our U-Net model:
In the following section, we investigate the hyperparameter ’s role in training the model. In particular, we consider , where for , and for , . We provide in Table 4 the results over the validation set of the five models trained with the different values of . As we can see, when the model is trained on the mean squared error alone, it focuses solely on minimizing the magnitude differences between the predicted and original ECG signals. Notably, while the results for RMSE and MAE remain consistent for the models with , the values for the PCC and DTW present higher variability, with an improvement of 4.65% in terms of PCC for the model with w.r.t. the model with and of 10.44% in terms of DTW. Interestingly, setting , resulted in the model performing well in terms of PCC but significantly worse across other metrics. This observation underscores the importance of employing a hybrid loss considering the ECG signal’s spatial and temporal aspects.
For our testing phase, we chose based on its performance during validation, showing promising results compared to and , without extensive exploration of other values. However, a more detailed examination revealed to potentially offer better outcomes. This finding suggests that while provided satisfactory results, a further optimization of could enhance model performance. In future work, we will investigate this aspect more thoroughly to identify the most effective parameter value.
@@
A.2. Loss evolution during training ()
Figure 7 illustrates the progression of the total loss (in blue), (in orange), and (in green) across 100 epochs. The total loss decreased, driven by the decline of the MSE and Pearson terms. The rapid reduction of the Pearson term during the early epochs indicates that the network quickly learns to align its predictions with the desired correlation structure. Meanwhile, the gradual decrease in MSE reflects ongoing improvements in reconstruction accuracy.
A.3. Impact of ECGrecover on signal quality assessment
SQI () | SQI % () | |||
Config. | Original | ECGrecover | Original | ECGrecover |
C1 | 0.520.09 | 0.520.08 | 623 | 317 |
C2 | 0.520.09 | 0.510.09 | 623 | 317 |
C3 | 0.520.09 | 0.510.09 | 623 | 316 |
C4 | 0.520.09 | 0.520.09 | 623 | 318 |
C5 | 0.520.09 | 0.480.10 | 623 | 318 |
C | 0.520.09 | 0.530.09 | 623 | 318 |
C | 0.520.09 | 0.530.08 | 623 | 213 |
C | 0.520.09 | 0.530.08 | 623 | 1636 |
C | 0.520.09 | 0.530.08 | 623 | 111 |
C | 0.520.09 | 0.530.09 | 623 | 624 |
C | 0.520.09 | 0.530.09 | 623 | 318 |
C | 0.520.09 | 0.530.09 | 623 | 215 |
C | 0.520.09 | 0.530.08 | 623 | 18 |
C | 0.520.09 | 0.530.08 | 623 | 215 |
C | 0.520.09 | 0.530.08 | 623 | 112 |
C | 0.520.09 | 0.530.08 | 623 | 19 |
C | 0.520.09 | 0.530.08 | 623 | 18 |
In this section, we examine the impact of ECGrecover’s reconstruction on ECG quality. ECG recordings are frequently susceptible to noise from various sources, including loose electrode contacts, patient movement, and muscle contractions. These factors can distort the waveform, complicating analysis and affecting diagnoses (granese2023negative). Therefore, we aim to verify that the proposed method does not degrade the quality of the reconstructed ECG compared to the original. It is important to note that our primary goal is not to improve ECG quality, as ECGrecover is not intended to be a denoising method.
For this set of analyses, we measured the ECG quality with standard signal quality indicators (SQI): the SQI (Makowski2021neurokit) and SQI (zhao2018sqi) methods, which are both already implemented in the Python library Neurokit (Makowski2021neurokit). SQI involves isolating the QRS complexes based on the R-peak locations. Each QRS complex is then standardized by subtracting the average QRS complex and dividing by the standard deviation. The average of these standardized QRS complexes is computed, and the results are re-scaled to a range between 0 and 1, with 1 indicating the highest quality. This method assumes that ECG signal segments exhibit a regular morphology, meaning that most QRS complexes are expected to be similar in shape (campero2023signal). SQI combines different quality indexes: power spectrum distribution of QRS wave, baseline relative power, and kurtosis of the ECG signal. The overall output is then used to classify the signal into one of three quality levels: excellent, barely acceptable, or unacceptable (zhao2018sqi). To numerically quantify the results regarding SQI, we calculated the percentage of ECGs classified as unacceptable over the total number of ECGs. In Table 5, we present the results of the two metrics computed for both the original ECGs in the test set of the Generepol dataset and those reconstructed with ECGrecover. Regarding SQI, the reconstructed signals maintain, on average, the same quality as the original signals, with fluctuations of 0.01 between the results of the two methods. In contrast, SQI shows a lower proportion of ECGs labelled as unacceptable with ECGrecover compared to the original dataset. This is evident in Figure 12, where we observe that ECGrecover attenuates overall noise in the signal more effectively than competitors in the C3 setting. Moreover, in the lead-reconstruction setting, we notice a positive influence from the other leads on the reconstructed one. This effect is illustrated in Figure 4 where the correction of the basal drift is performed for signal in the C setting.
Finally, we would like to emphasize that no additional noise was added to the signals in our comparison, as the proposed method is not intended to be a denoising technique. Additionally, it is important to note that the concept of ECG quality is not perfectly defined and may vary depending on the measurement criteria. Nevertheless, these results support the usability of ECGrecover as a reconstruction method in real-life settings.
A.4. Random masking
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | |
@@
An illustration of an ECG subjected to such random masking, C, is depicted in Figure 8. For this process, a random start and end point between 0 and the total length of the signal were selected for each ECG lead. The segments outside the range between these two points were masked.
@@
The results, detailed in Table 6, indicate that ECGrecover consistently outperforms its competitors across various metrics, even in this challenging scenario. For instance, in the most difficult case concerning lead III, ECGrecover not only improves upon the best competitor, Pix2Pix, by 51% in terms of PCC but also surpasses it by approximately 77.78% in terms of DTW. Furthermore, ECGrecover demonstrates stable performance across different leads, contrasting with the variability seen in competitors like EKGAN and Pix2Pix, which show wide fluctuations in PCC and DTW scores.
A.5. Demographic analysis and impact on model performance
A.5.1. Demographic in Generepol
Metric | Age | Height (cm) | Weight (kg) |
Training set: 800 patients (39% F and 61% M) | |||
2810 | 170.58.8 | 65.610.1 | |
24 | 170 | 65 | |
18 | 150 | 45 | |
60 | 197 | 103 | |
Validation set: 100 patients (40% F and 60% M) | |||
31.512 | 170.19.6 | 66.89.8 | |
25 | 169 | 66 | |
18 | 152 | 46 | |
60 | 191 | 95 | |
Testing set: 89 patients (30% F and 60% M) | |||
258 | 170.47.9 | 64.410 | |
22 | 168 | 62 | |
18 | 151 | 49 | |
57 | 188 | 102 |
We provide in Table 7 the summary of demographic information regarding the Generepol dataset. The study included 989 patients with no known heart problems. For a deeper explanation, we refer to (salem2017genome).
A.5.2. Performance of ECGrecover across demographic groups.
We consider the linear mixed model as our statistical approach to understanding how various predictors (e.g., height, weight, and sex) influence the dependent variables PCC, RMSE, and DTW. In this case, the random effect is the intercept for the ECG. The model was fitted using restricted maximum likelihood and Satterthwaite’s method for t-tests.
In Figures 9(c), 10(c) and 11(c), we display the statistical analysis results. The x-axis represents the estimated effect size, which measures the relationship between each predictor and the outcome metric. This effect size indicates the expected change in the outcome for a one-unit change in the predictor while keeping all other predictors constant. The distance of each dot from the vertical zero line on the x-axis reflects the magnitude of the effect size, with greater distances signifying stronger effects. Additionally, asterisks next to the dots denote the statistical significance of the estimates: one asterisk () for , two asterisks () for , and three asterisks () for . These asterisks highlight which predictors have a statistically significant impact on the outcome metrics.
The plots show that ’weight’ and ’height’ do not affect performance metrics across all methods, as their effect size is zero. The same holds for the predictor ’age’ concerning the PCC metric, where for ECGrecover, the effect size is very small (¡1e-3), cf. Figure 9(c).
Conversely, the estimates for ’age’ on RMSE and DTW are very small (close to zero) but statistically significant, indicating a very small but real effect of ’age’ on these variables that is not due to random variation (cf. Figures 10(c) and 11(c)). This observation holds for ECGrecover and EKGAN but not for Pix2Pix.
Furthermore, we observed that the time of sotalol administration (3 hours after drug intake) significantly impacts all performance metrics, with a higher effect size (e.g., greater than a 0.01 decrease in PCC). We hypothesize that the drug’s footprint or ’age’ at a lower level makes it more challenging for all the methods to reconstruct the original ECG from the masked signal.
Finally, the ’sex’ predictor shows a slightly negative effect for Pix2Pix across all metrics, indicating poorer performance in male participants.
A.6. ECGrecover on PTB-XL (wagner2020ptb)
@@
We present the evaluation of the proposed ECGrecover on PTB-XL (wagner2020ptb). Specifically, the dataset contains 21799 recordings from 18869 patients of 10-second 12- lead ECGs sampled at 500Hz labeled as normal, myocardial infarction, ST/T change, conduction disturbance and hypertrophy. In Figure 14, we show the summary of the results. In particular, we followed the same experimental setting as the one already described for the Generepol dataset in Section 3.1.To get an overall view, we also include the same representation of the results for the Generepol dataset in the figure.
A.7. Experimental environment and computational costs
In Table 8, we summarise the computational time for training and testing ECGrecover and the competitors.
ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | |
Training time | |||
Time per ECG | 4ms | 4ms | 3ms |
N. Parameters | 6.147.982 | 26.122.880 | 13.915.393 |
GPU memory quota | 0.6 GB | 11.5 GB | 11.8 GB |
Inference time | |||
Time per ECG | 2ms | 1ms | 1ms |
Appendix B Comprehensive set of numerical results
B.1. Additional plots
@@
We include additional plots in Figure 12 and Figure 13, which could not be accommodated in the main paper due to space limitations. Specifically, in Figure 12, we illustrate a lead II signal reconstruction from input masks C3 and C5 across all methodologies evaluated in our study. Figure 13 presents the ECG waveforms alongside their labelled wave coordinates (peaks) to enhance comprehension of the findings discussed in Section 3.2.2.
B.2. Additional tables
@@
This section includes the comprehensive numerical results of the evaluation. Specifically, Tables 9, 11 and 10 present the results for the PCC metric; Tables 12, 14 and 13 the results for RMSE metric; Tables 15, 17 and 16 the results for MAE metric; Tables 18, 20 and 19 the results for DTW metric.
@@
Finally, in Tables 21, 23 and 22 we include the results on the QT-segment preservation w.r.t. all the possible input masks. Additionally, we expanded the analysis on the QRS complex (see Figure 13) in Tables 24, 26 and 25. Moreover, in Tables 27, 29 and 28, we show the percentage of R peaks detected (%-R) in the reconstructed signals.
We highlight that the content of the tables has been analyzed in the paper through various plots and summaries of the results. However, for transparency, we also consider providing the full set of numerical results important.
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |